OpenStudy (anonymous):

Given the parent functions f(x) = 5x-1 and g(x) = 3^x -9 what is gx-fx

OpenStudy (loser66):

just subtract them,

hartnn (hartnn):

like this : \(\large g(x) - f(x) = 3^x-9 - (5x-1)\) simplify

OpenStudy (anonymous):

i got it :D

OpenStudy (anonymous):

Can you help with another

hartnn (hartnn):

sure :)

OpenStudy (anonymous):

given the parent functions f(x) = log3 (5x-5) and gx = log 3 (x-1) what is f(x) - g(x)

hartnn (hartnn):

did you try to subtract them ? then use the log property that \(\large \log_ab -\log_a c = \log_a (\dfrac{b}{c})\)

OpenStudy (anonymous):

no..so I should do the numbers first before log?

hartnn (hartnn):

no, i meant this : \(\large f(x)-g(x) = \log_3 (5x-5)- \log_3 (x-1 ) \\ \large = \dfrac{\log_3 (5x-5)}{\log_3(x-1)}\) got this ?

OpenStudy (anonymous):

I never really understood log at alll!

hartnn (hartnn):

sorry, that should be \(\large f(x)-g(x) = \log_3 (5x-5)- \log_3 (x-1 ) \\ \large = \log_3 \dfrac{ (5x-5)}{(x-1)}\) and logs are not difficult. here, i used the property that , differences of log equals log of ratio

OpenStudy (anonymous):

Oh ok I see what you did there! What should be the next step?

hartnn (hartnn):

factor out 5 from 5x-5

hartnn (hartnn):

5(x-1) and not that x-1 gets cancelled from numerator and denominator :)

OpenStudy (anonymous):

yeah so x-1 will stay and you have 5(x-1)

hartnn (hartnn):